Find the critical value SELECT ALL APPLICABLE CHOICES A) 0.412 B) 0.922 C) 16.75 D) 1.078 corresponding to a sample size of 6 and a confidence level of 99.0 percent Determine the corresponding values for a 92.5 percent confidence interval SELECT ALL APPLICABLE CHOICES B) C) 0.2 in/2 = ±-1.85 0.11 92.5% F) None of These Of 649 adults selected randomly from one town, 54 of them smoke. Construct a 92.5 percent n interval for the true percentage of all adults in the town that smoke SELECT ALL APPLICABLE CHOICES A) 0.0639 <p< 0.10251 0.0139< p<0.13251 C) None of These
Answer #1

1)

Critical value of chi square distribution for n-1=5 df and 99 % CI  
Lower critical value χ2L= 0.412

option A is correct

2)

since it is 2 tailed test:

2a/2 = +-1.78

3)

sample success x = 54
sample size          n= 649
sample proportion p̂ =x/n= 0.0832
std error se= √(p*(1-p)/n) = 0.0108
for 92.5 % CI value of z= 1.780
margin of error E=z*std error   = 0.0193
lower bound=p̂ -E                       = 0.0639
Upper bound=p̂ +E                     = 0.1025

0.0639 <p <0.10251